OpenStudy (anonymous):

14. What is the sum of the arithmetic series 18 (on top) series symbol t=1 on bottom then (5t-4) a. 783 b1566 c819 d86

OpenStudy (anonymous):

@Opcode

OpenStudy (opcode):

Summation right? So you mean this: \[\sum_{t=1}^{18}(5t-4)\]

OpenStudy (anonymous):

YES!!!

OpenStudy (opcode):

Have you learned basic summation? What do you not understand?

OpenStudy (anonymous):

I have I just drew a complete blank.

OpenStudy (opcode):

Hmm, okay do you want to attempt the problem on your own, and then want me to check your answer? I'm sure you can do it the problem isn't too hard, if you don't know where to start just ask me. :-)

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

first term a=5*1-4=1 last term l=5*18-4=90-4=86 n=18 \[sn=\frac{ n }{ 2 }\left( a+l \right)\] calculate s18

OpenStudy (opcode):

No, it is not C or D. Summation works like: \[\sum_{t \mathop =m}^n a_t = a_m + a_{m+1} + a_{m+2} +\cdots+ a_{n-1} + a_n\]

OpenStudy (opcode):

T represents the index of the summation. M represents the lower limit of the summation. N represents the upper limit of the summation. \[\sum_{t=1}^{18}(5t-4)\] Think of it like 18 is the number we want to go to, and that t = 1 and 1 being the start. Here is a simple example to get you started: \[\sum_{n=2}^{5}(n) = 2 + 3 + 4 + 5 = 14\] Do you see how that works?

OpenStudy (anonymous):

Yes I understand that

OpenStudy (anonymous):

Oh I understand thank you

OpenStudy (opcode):

Okay here is one that is a little bit more complex: \[\sum_{x = 1}^{5}(x^2 + 2)\] Now do you see how to do that? Here is a hint: \[\sum_{x = 1}^{5}(x^2 + 2) = (1^2 + 2) + (2^2 + 2) + (3^2 + 2) + (4^2 + 2) + (5^2 + 2)\] See how x = 1 is the number we start at and the 5 on top of the capital sigma the number we go to?

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

I got it right(: Thank you so much

OpenStudy (opcode):

Okay, well I might as well right out the answer just for others who may read this: \[\sum_{t=1}^{18}(5t-4) = (5 \times 1 - 4) + (5 \times 2 - 4) + (5 \times 3 - 4) \dots = 783\] If you don't understand just say. :-)

OpenStudy (anonymous):

Thank you bunches

OpenStudy (opcode):

No problem, glad you understand. :D